# Enhancing HNSW Index for Real-Time Updates: Addressing Unreachable Points and Performance Degradation

### v1

@@ Abstract
The approximate nearest neighbor search (ANNS) is a fundamental and essential component in information retrieval, with graph-based methodologies demonstrating superior performance compared to alternative approaches. Extensive research efforts have been dedicated to improving search efficiency by developing various graph-based indices, such as HNSW (Hierarchical Navigable Small World). However, the performance of HNSW and most graph-based indices becomes unacceptable when faced with a large number of real-time deletions, insertions, and updates. Furthermore, during update operations, HNSW can result in some data points becoming unreachable, a situation we refer to as the ‘unreachable points phenomenon’. This phenomenon could significantly affect the search accuracy of the graph in certain situations.
To address these issues, we present efficient measures to overcome the shortcomings of HNSW, specifically addressing poor performance over long periods of delete and update operations and resolving the issues caused by the unreachable points phenomenon. Our proposed MN-RU algorithm effectively improves update efficiency and suppresses the growth rate of unreachable points, ensuring better overall performance and maintaining the integrity of the graph. Our results demonstrate that our methods outperform existing approaches. Furthermore, since our methods are based on HNSW, they can be easily integrated with existing indices widely used in the industrial field, making them practical for future real-world applications. Code is available at https://github.com/xwt1/ICPADS-MN-RU.git
@@
Information retrieval, Incremental Update, Freshness, Graph-Based Index
@@ II-B HNSW
    Collect Neighbors: The algorithm gathers the one-hop and two-hop neighbors of the deleted point. For each one-hop neighbor $v_{j}$, the one-hop and two-hop neighbors and the newly inserted point are considered potential candidates.
@@ III-A Causes of Unreachable Points
In HNSW, the occurrence of some points having an in-degree of zero after the replaced_update operation is attributable to the specific repair connection strategy employed by the algorithm. When performing a replaced_update, HNSW first collects the one-hop neighbors of the points to be deleted. Then, it sets the two-hop neighbors combined with the one-hop neighbors, including the newly inserted point, as new candidate neighbors for each one-hop neighbor and prunes this candidate set to a suitable size. The edges from the one-hop neighbors to their original neighbors might be deleted during these operations. As a result, these original neighbors may end up with no in-edges after the replaced_update operation. Consequently, these points will not be found during future search processes unless they are the entry points of the HNSW.
@@ III-B Demonstration on Specific Datasets
Figure 3 (a) illustrates that the number of unreachable points in HNSW consistently increases with continuous deletion and reinsertion. After approximately 3000 iterations, the number of unreachable points reaches between 3% and 4% of the original dataset. This number rises with additional iterations, indicating that points already in the unreachable set will not be searched in future search processes.
Figure 3 (b) shows that recall decreases as the number of iterations increases. With the same configuration, recall declines by about 3% due to the increasing number of unreachable points over iterations. This decline becomes more severe with more iterations, leading to a gradual reduction in the accuracy of the final search. Furthermore, this decrease in accuracy caused by the unreachable points phenomenon cannot be mitigated by adjusting HNSW searching parameters, such as increasing the ef_ parameter.
@@ IV-A Dual Indexes Design
Diverging from conventional methodologies reliant on a singular HNSW index, we introduce an additional index designed explicitly for managing unreachable points within the HNSW indexing system, referred to as the HNSW Backup Index, as illustrated in Fig. 4. The inception of this algorithm is rooted in minimizing the computational burden and service disruptions associated with recurrent reindexing processes. The procedure for creating the Backup Index involves several steps: First, the HNSW K-NN search function is applied to identify the point set $F$ that is closest to the query point $q$ from the current data set $P$. We set $K$ to $|P|$ to retrieve all the points available in the index. Next, by removing $F$ from the set $P$, the remaining set of points $U$, which includes the points not found, is identified and extracted. Subsequently, a new HNSW backup index is constructed for these unfound point sets $U$. This process leverages the multi-level proximity search feature of HNSW to efficiently manage unreachable points, ensuring that the structure and query performance of the original index remains unaffected while reducing the computational overhead and service interruptions typically associated with reconstruction.
@@ IV-B Dual Index Search
Algorithm 1 outlines the search process for dualSearch. Initially, it performs a K-NN search on the primary index ($HNSW_{main}$) using the query point $q$ to retrieve the nearest neighbors $k$ in the set of results $R_{main}$. Concurrently, a similar K-NN search is executed on the backup index ($HNSW_{backup}$), dedicated to managing unreachable points, yielding the results set $R_{backup}$. Subsequently, the results of both searches are merged into a unified set $R_{combined}$, which is then sorted according to the distance from point to query to prioritize closer points. Ultimately, the algorithm delivers the top $k$ points from this ordered set, ensuring a precise and effective query result that takes into account both the accessible and the unreachable points.
@@ IV-C Index Maintenance
As mentioned earlier, the original HNSW replaced_update algorithm needs to consider the neighbors of the deleted point(namely one-hop neighbors) and the neighbors of these neighbors (namely two-hop neighbors) as potential new neighbors to one-hop neighbors during the update process. For each neighbor, the algorithm has to reselect its neighbors from a candidate set of size $M^{2}$, including its neighbors and the two-hop neighbors. Here, $M$ is a parameter in HNSW that defines the maximum deviation of data points within a specific layer. The time complexity of this operation is $O(M^{3})$ per layer, which is significantly time-consuming.
@@
To address this complexity, we propose a novel algorithm, Mutual Neighbor Replaced Update (MN-RU), which focuses on restoring the neighbors’ connectivity of the deleted point. This process ensures index connectivity and enhances search efficiency. In the context of a deleted point’s neighbors, it is noted that most two-hop neighbors of the deleted point may not be the neighbors of its one-hop neighbors. Considering all two-hop neighbors as candidates is inefficient. Instead, optimizing update efficiency and maintaining connectivity can be achieved by combining the original neighbors of each deleted point’s neighbor with the neighbors of the deleted point to form the candidate set. Algorithm 2, contrary to the original method, exclusively updates neighbors of the deleted point that include the deleted point in their neighbor list. For every such point, the candidate neighbors comprise its original neighbors along with the neighbors of the deleted point. This operation exhibits a time complexity of $O(M^{2})$ per layer since each update’s set of candidate neighboring is on scale $M$. As a result, our algorithm demonstrates superior efficiency compared to the original approach. Figure 5 serves as an illustrative example of MN-RU.
The complete process of Index Maintenance is detailed below:
Deletion. For deletions, as illustrated in Figure 4, when a point in the index is deleted, our method appends it to a list, deletedList, similar to the process in the original HNSW. The purpose of utilizing the deletedList is to manage deleted points without immediate removal efficiently, thus preserving the graph’s structural integrity during updates. This deferred deletion approach reduces re-indexing overhead and allows for efficient reuse of memory space when new points are inserted, optimizing memory usage, and preventing excessive index expansion.
@@
Insertion. Algorithm 2 and Algorithm 3 presents a method for updating the HNSW graph when inserting a new point to replace a deleted point. Algorithm 2 is first applied to repair the graph’s connectivity, followed by Algorithm 3, which adds the new point to the index. Unlike the conventional approach of adding new points, the Algorithm 3 enables the new point to inherit the layer level of the deleted point. Subsequent to this inheritance, the new point undergoes insertion using the standard HNSW insert process. Initially, the algorithm identifies the maximum layer of the deleted point and top layer of the existing HNSW graph and establishes an initial entry point for search. Progressing from the top layer to the layer above the maximum layer of the deleted point, it iteratively searches for the nearest point to the new data point, adjusting the entry point after each iteration. Subsequently, from the maximum layer of the deleted point to the lowest layer, it conducts searches for neighboring points at each layer, selects neighbors for the new point, and establishes connections between the new point and its neighbors, updating the entry point iteratively. This meticulous process ensures the seamless integration of the new element into the HNSW structure.
@@ V-B Baselines
Given the limited support for online update operations in the majority of memory-based ANNS methods [7, 8, 32], our primary comparison centers on contrasting our methods with HNSW’s replaced_update algorithm. Derived mainly from Algorithm 2, our methods emphasize the requirement for mutually connected points with deleted points for neighbor reselection, excluding MN-THN-RU, which will be discussed later. These methods are collectively termed Mutual Neighbor Replaced Update (MN-RU) with distinctive suffixes. Hereafter, we introduce both the baseline methods and our strategy:
@@
  • Mutual Neighbor replaced_update $\beta$(MN-RU $\beta$): Similar to the previous method, MN-RU $\beta$ selects the point set $P$ associated with the deleted points to reselect neighbors. For each point $v$ in $P$, it includes the neighbors of the deleted points and $v$’s neighbors as the candidate new neighbor set $C$. The $\alpha$-RNG pruning strategy [27] is applied with $\alpha=1$.
@@
  • Mutual Neighbor And Two Hop Neighbor replaced_update (MN-THN-RU): A variant of MN-RU $\gamma$, MN-THN-RU includes points mutually connected with deleted points and neighbors of neighbors of deleted points that have a connection to the deleted point.
@@
V-C Experimental Scenarios
This section compares our methods against the HNSW replaced_update algorithm in the following three scenarios.
@@ V-C Experimental Scenarios
  • New_Data: Focusing on the Sift2M dataset, we initialize the index with the first million data points and undertake 10 iterations. Each iteration involves the deletion of a portion from the first million and the insertion of a component from the second million, aiming to assess the performance and adaptability of the index with the continuous introduction of new data.
@@
This comparison relies on two key metrics: update time efficiency and growth of unreachable points. A method with superior efficiency exhibits reduced update times, while a method demonstrating excellence showcases minimal growth in unreachable points per iteration. We evaluated our methods against the HNSW replaced update algorithm in three scenarios, focusing on update time and the number of unreachable points.
@@ V-D Update time efficiency
Therefore, we investigated the search performance of indices built on the Gist and ImageNet datasets after updates to assess how unreachable points affect recall. The original Gist query set has only 1,000 queries and the ImageNet query set has 200, making it difficult to cover the entire dataset, especially for small values of $K$ ($K\leq 100$). This limits our ability to observe the impact of unreachable point growth on search accuracy.
@@ VI Conclusion
This paper addresses HNSW limitations in real-time deletions and insertions, leading to the ‘unreachable points phenomenon’ and reduced efficiency. Our proposed MN-RU and MN-THN-RU algorithms with the backup index method effectively mitigate these issues by enhancing the efficiency of mixed operations and maintaining better graph connectivity. Extensive experimental validation demonstrates that our methods outperform the native HNSW strategy, significantly reducing the number of unreachable points and improving the update speed across various datasets. These improvements make our approach highly practical for real-world applications requiring dynamic real-time updates and high search accuracy.

### v2

@@ Abstract
The approximate nearest neighbor search (ANNS) is a fundamental and essential component in data mining and information retrieval, with graph-based methodologies demonstrating superior performance compared to alternative approaches. Extensive research efforts have been dedicated to improving search efficiency by developing various graph-based indices, such as HNSW (Hierarchical Navigable Small World). However, the performance of HNSW and most graph-based indices become unacceptable when faced with a large number of real-time deletions, insertions, and updates. Furthermore, during update operations, HNSW can result in some data points becoming unreachable, a situation we refer to as the ‘unreachable points phenomenon’. This phenomenon could significantly affect the search accuracy of the graph in certain situations.
To address these issues, we present efficient measures to overcome the shortcomings of HNSW, specifically addressing poor performance over long periods of delete and update operations and resolving the issues caused by the unreachable points phenomenon. Our proposed MN-RU algorithm effectively improves update efficiency and suppresses the growth rate of unreachable points, ensuring better overall performance and maintaining the integrity of the graph. Our results demonstrate that our methods outperform existing approaches. Furthermore, since our methods are based on HNSW, they can be easily integrated with existing indices widely used in the industrial field, making them practical for future real-world applications. Code is available at https://github.com/xwt1/MN-RU.git
@@
Data Mining, Information Retrieval, Incremental Update, Freshness, Graph-Based Index
@@ II-B HNSW
    Collect Neighbors: The algorithm gathers the one-hop and two-hop neighbors of the deleted point. For each one-hop neighbor $v_{j}$, the one-hop and two-hop neighbors and the newly inserted point are considered potential candidates new neighbors for $v_{j}$.
@@ III-A Causes of Unreachable Points
In HNSW, the occurrence of some points having an in-degree of zero after the replaced_update operation is attribute to the specific repair connection strategy employed by the algorithm. During a replaced_update in HNSW, the algorithm first gathers the neighbors of the deleted point, referred to as $T$. It then combines $T$ with their own neighbors and the newly inserted point to form a new candidate neighbor set for each member of $T$. Finally, new neighbors are selected from the candidate set for each member of $T$. The edges from members of $T$ to their original neighbors might be deleted during these operations. As a result, these original neighbors may end up with no in-edges after the replaced_update operation. Consequently, these points will not be found during future search processes unless they are the entry points of the HNSW.
@@ III-B Demonstration on Specific Datasets
Figure 3 (a) illustrates that the number of unreachable points in HNSW consistently increases with continuous deletion and reinsertion. After approximately 3000 iterations, the number of unreachable points reaches between 3% and 4% of the original dataset. This number rises with additional iterations, indicating that points already in the unreachable points set will not be searched in future search processes.
Figure 3 (b) shows that recall decreases as the number of iterations increase. With the same configuration, recall declines by about 3% due to the increasing number of unreachable points over iterations. This decline becomes more severe with more iterations, leading to a gradual reduction in the accuracy of the final search. Furthermore, this decrease in accuracy caused by the unreachable points phenomenon cannot be mitigated by adjusting HNSW searching parameters, such as increasing the ef_ parameter.
@@ IV-A Dual Indexes Design
Diverging from conventional methodologies reliant on a singular HNSW index, we introduce an additional index designed explicitly for managing unreachable points within the HNSW indexing system, referred to as the HNSW Backup Index, as illustrated in Figure 4. The inception of this algorithm is rooted in minimizing the computational burden and service disruptions associated with recurrent reindexing processes. The procedure for creating the Backup Index involves several steps: First, the HNSW K-NN search function is applied to identify the point set $F$ that is closest to the query point $q$ from the current data set $P$. We set $K$ to $|P|$ to retrieve all the points available in the index. Next, by removing $F$ from the set $P$, the remaining set of points $U$, which includes the points not found, is identified and extracted. Subsequently, a new HNSW backup index is constructed for these unfound point sets $U$. This process leverages the multi-level proximity search feature of HNSW to efficiently manage unreachable points, ensuring that the structure and query performance of the original index remains unaffected while reducing the computational overhead and service interruptions typically associated with reconstruction.
@@ IV-B Dual Index Search
Algorithm 1 outlines the search process for dualSearch. Initially, it performs a K-NN search on the primary index ($HNSW_{main}$) using the query point $q$ to retrieve $k$ nearest neighbors in the set of results $R_{main}$. Concurrently, a similar K-NN search is executed on the backup index ($HNSW_{backup}$), dedicated to managing unreachable points, yielding the results set $R_{backup}$. Subsequently, the results of both searches are merged into a unified set $R_{combined}$, which is then sorted according to the distance from point to query to prioritize closer points. Ultimately, the algorithm delivers the top $k$ points from this ordered set, ensuring a precise and effective query result that takes into account both the accessible and the unreachable points.
@@ IV-C Index Maintenance
As mentioned earlier, the original HNSW replaced_update algorithm needs to consider the neighbors of the deleted point(namely one-hop neighbors) and the neighbors of these neighbors (namely two-hop neighbors) as potential new neighbors to one-hop neighbors during the update process. For each one-hop neighbors, the algorithm has to reselect its neighbors from a candidate set of size $M^{2}$, including one-hop neighbors and the two-hop neighbors. Here, $M$ is a parameter in HNSW that defines the maximum out-degree of a data point within a specific layer. The time complexity of this operation is $O(M^{3})$ per layer, which is significantly time-consuming.
@@
For selecting new neighbors for one-hop neighbors, considering all two-hop neighbors as candidates is inefficient. The reason is as follows: Before update, based on HNSW’s edge selection strategy, if a one-hop neighbor does not have an edge to a particular two-hop neighbor, it means HNSW deemed this edge suboptimal. Since the size of two-hop neighbors is $M^{2}$ and each one-hop neighbor originally has $M$ neighbors, many two-hop neighbors are not part of its original neighbor set. Therefore, establishing edges from one-hop neighbor to such two-hop neighbors was deemed suboptimal by HNSW before update and is likely to remain so after update, as the index structure does not undergo significant changes.
To address this complexity, we propose a novel algorithm(Algorithm 2) which focuses on restoring the connectivity of one-hop neighbors of the deleted point. Contrary to original method, Algorithm 2 exclusively updates neighbors of one-hop neighbors that include the deleted point in their original neighbor list. For every such point, the candidate new neighbors comprise its original neighbors along with the neighbors of the deleted point. Due to the previously described reasons, two-hop neighbors are not necessarily included in the new neighbor candidate set, ensuring a more efficient and focused neighbor update process. This operation exhibits a time complexity of $O(M^{2})$ per layer since for each one-hop neighbors, the algorithm has to reselect its neighbors from a candidate set of size $M$. Therefore, our algorithm demonstrates superior efficiency compared to the original approach. Figure 5 serves as an illustrative example of Algorithm 2. The complete process of Index Maintenance is detailed below:
Deletion. For deletions, as illustrated in Figure 4, when a point in the index is deleted, our method appends it to a list, deletedList, similar to the process in the original HNSW. The purpose of utilizing the deletedList is to manage deleted points without immediate removal efficiently, thus preserving the graph’s structural integrity during updates.
@@
Insertion. Algorithms 2 and 3, collectively called Mutual Neighbor Replaced Update (MN-RU), update the HNSW index when a new point replaces a deleted one. First, Algorithm 2 repairs the graph’s connectivity, then Algorithm 3 inserts the new point to the index. Unlike the conventional approach of adding new points, the Algorithm 3 enables the new point to inherit the layer level of the deleted point. Subsequent to this inheritance, the new point undergoes insertion using the standard HNSW insert process. Algorithm 3 starts by identifying the top layer of the HNSW index and the maximum layer of the deleted point, establishing an initial entry point for the search. From the top layer to the layer above the deleted point’s maximum layer, it iteratively searches for the nearest point as the entry point to the new data point, updating the entry point each time. Then, from the maximum layer of the deleted point to the lowest layer, it searches for neighbors, selects neighbors for the new point, and establishes connections. MN-RU ensures the seamless integration of the new element into the HNSW structure.
@@ V-B Baselines
Given that most memory-based ANNS methods [7, 8, 32] do not support real-time update operations, our primary comparison centers on contrasting our methods with HNSW’s replaced_update algorithm. Our methods, based mainly on Algorithm 2, focus on reselecting neighbors only for points mutually connected with the deleted points, except for MN-THN-RU, which will be discussed later. These methods are collectively termed Mutual Neighbor Replaced Update (MN-RU) with distinctive suffixes. Hereafter, we introduce both the baseline methods and our strategy:
@@
  • Mutual Neighbor replaced_update $\beta$(MN-RU $\beta$): Similar to the previous method, MN-RU $\beta$ selects the point set $P$ with mutual connections to the deleted points for neighbor reselection. For each point $v$ in $P$, it includes the neighbors of the deleted points and $v$’s original neighbors as the candidate new neighbor set $C$ for $v$. The $\alpha$-RNG pruning strategy [27] is applied with $\alpha=1$.
@@
  • Mutual Neighbor And Two Hop Neighbor replaced_update (MN-THN-RU): A variant of MN-RU $\gamma$, MN-THN-RU reselect the neighbors for points that mutually connected with deleted points and neighbors of neighbors of deleted points that have a connection to the deleted point.
@@
V-C Experimental Scenarios And Metrics
We compared our methods against the HNSW replaced_update algorithm in the following three scenarios.
@@ V-C Experimental Scenarios And Metrics
  • New_Data: Focusing on the Sift2M dataset, we initialize the index with the first million data points and conduct 10 iterations. Each iteration deletes 100,000 points from the first million and inserts 100,000 points from the second million. This process aims to assess the index’s performance and adaptability with continuous data introduction. The final index consists of the second million data points.
@@
This comparison relies on two key metrics: update time efficiency and growth of unreachable points. A method with superior efficiency exhibits reduced update times, while a method demonstrating excellence showcases minimal growth in unreachable points per iteration. We evaluated our methods against the HNSW replaced_update algorithm in three scenarios, focusing on update time and the number of unreachable points.
@@ V-E Unreachable Points Growth
Therefore, we investigated the search performance of indices built on Gist and ImageNet datasets under full_coverage and random scenarios after updates to assess how unreachable points affect recall. The original Gist query set has only 1,000 queries and the ImageNet query set has 200, making it difficult to cover the entire dataset, especially for small values of $K$ ($K\leq 100$). This limits our ability to observe the impact of unreachable point growth on search accuracy.
@@ VI Conclusion
This paper addresses HNSW replaced_update limitations in real-time deletions and insertions, leading to the ‘unreachable points phenomenon’ and reduced efficiency. Our proposed MN-THN-RU and MN-RU $\gamma$ algorithms with the backup index method effectively mitigate these issues by enhancing the efficiency of mixed operations and maintaining better graph connectivity. Extensive experimental validation demonstrates that our methods outperform the native HNSW strategy, significantly reducing the number of unreachable points and improving the update speed across various datasets. These improvements make our approach highly practical for real-world applications requiring dynamic real-time updates and high search accuracy.